# Dense-SfM: Structure from Motion with Dense Consistent Matching

### v1

@@ 4.1 Datasets
We use the ETH3D [53] and the Texture-Poor SfM datasets [24], each presenting unique challenges across different types of scenes. The ETH3D dataset includes 25 indoor and outdoor scenes with sparsely captured, high-resolution images, and ground truth poses calibrated using Lidar. The Texture-Poor SfM dataset focuses on object-centric scenes with low texture, comprising subsampled image sets. To obtain dense matching from image pairs, we resize images to $1162\times 768$ on the ETH3D dataset, while keeping the original size of $840\times 840$ on the Texture-Poor SfM dataset.
@@ 4.2 Implementation Details
For efficiency, we apply non-max suppression [8] with a radius of 4 pixels for ETH3D and 2 pixels for Texture-Poor SfM. For mutual verification, we filter out matches with a distance exceeding $\epsilon_{\textit{p}}=3$px in Eq. 2.
To determine the visibility of Gaussians (SfM points) from a camera pose, we set the visibility threshold $\epsilon_{\textit{v}}$ to 0.5 in Eq. 3. Details on parameters and training implementation of Gaussian Splatting are provided in the supplementary material.
We use S2Dnet [17] as a CNN backbone for our refinement architecture, and adopt the multi-view transformer architecture from [24]. To ensure a fair comparison, we apply the same refinement parameters as in [24], including a local patch size ($p=15$) from the feature maps ($\mathbf{F}_{R}$ and $\mathbf{F}_{Q}$) obtained from the CNN backbone, and a search boundary for reference view’s refined keypoints set to $w=7$. The number of distribution anchors $C^{2}$ is set to $7^{2}$. For the CNN decoder in Fig. 4, we use a series of residual blocks [22] and channel attention blocks [71]. During bundle adjustment, matches with reprojection errors exceeding $\epsilon_{\textit{f}}=3$px are discarded. The total SfM refinement process is iterated twice.
@@
4.3 3D Triangulation
Evaluation.
We use the ETH3D training set, comprising 13 indoor and outdoor scenes with millimeter-accurate dense point clouds as ground truth. Following the protocol in [11, 30], we triangulate 3D points using fixed camera poses and intrinsics, then evaluate the resulting point clouds for accuracy and completeness using the ETH3D benchmark [53]. The results are reported at varying distance thresholds (1cm, 2cm, 5cm), averaged across all scenes. Consistent with the benchmark setup, accuracy and completeness are measured by the proportion of triangulated points that fall within the specified distances from the ground truth points.
@@ Results.
The results are shown in Tab. 1. While there is a trade-off between accuracy and completeness, our Dense-SfM framework achieves superior performances on both metrics. Compared to the state-of-the-art baseline LoFTR + DFSfM, our pipeline with RoMa achieves higher accuracy and completeness, thanks to employing dense matching and our proposed pipeline. Also, compared to the RoMa + DFSfM, where we apply quantization on dense matching and refine the tracks as in [24], our method achieves higher performance in terms of both accuracy and completeness, highlighting the effectiveness of our SfM framework. More results are reported in the supplementary material.
Type Method ETH3D Dataset Texture-Poor SfM Dataset
AUC@1°$\degree$ AUC@3°$\degree$ AUC@5°$\degree$ AUC@3°$\degree$ AUC@5°$\degree$ AUC@10°$\degree$
Detector-Based COLMAP (SIFT+NN) [51] 26.71 38.86 42.14 2.87 3.85 4.95
SIFT [31] + NN + PixSfM [30] 26.94 39.01 42.19 3.13 4.08 5.09
D2Net [10] + NN + PixSfM 34.50 49.77 53.58 1.54 2.63 4.54
R2D2 [44] + NN + PixSfM 43.58 62.09 66.89 3.79 5.51 7.84
SP [8] + SG [49] + PixSfM 50.82 68.52 72.86 14.00 19.23 24.55
Semi-Dense Matching LoFTR [57] + PixSfM 54.35 73.97 78.86 20.66 30.49 42.01
LoFTR + DFSfM [24] 59.12 75.59 79.53 26.07 35.77 45.43
AspanTrans. [5] + DFSfM 57.23 73.71 77.70 25.78 35.69 45.64
MatchFormer [68] + DFSfM 56.70 73.00 76.84 26.90 37.57 48.55
Dense Matching RoMa + DFSfM 57.97 77.02 81.59 34.28 50.01 69.01
DKM + Ours 59.04 77.73 82.20 41.19 57.26 72.95
RoMa + Ours 60.92 78.41 82.63 43.37 60.68 77.87
@@ 4.4 Multi-View Camera Pose Estimation
For the evaluation, we use a subset of ETH3D training and test dataset, which includes 22 indoor and outdoor scenes, and the Texture-Poor SfM dataset. We evaluate the accuracy of the estimated multi-view poses using the AUC of pose error at various thresholds, following the evaluation protocol used in [30, 24].
@@ Results.
Figure 6: Qualitative Results on Texture-Poor SfM dataset. Our method with the dense matcher RoMa is qualitatively compared with the baseline LoFTR + DFSfM. The red cameras are ground truth poses while the blue cameras are recovered poses.
@@ 4.5 Ablation Studies
We conduct several ablation studies to evaluate the key components of our pipeline, including the matching design used as input for SfM and the architectural choices of the refinement module, on the ETH3D dataset using triangulation metrics.
@@
For a comparison with [24], which applies quantization to matched keypoints to address the fragmented track problem, we replaced our GS track extension with their quantization approach, using a rounding ratio of $r=4$. Results indicate that our GS track extension not only maintains completeness by avoiding the limitations of quantization but also achieves superior accuracy.
@@
We first compare our refinement module to the multi-view refinement module in [24], by substituting our refinement module with it. The result shows that our refinement module consistently outperforms the existing approach, achieving significantly higher accuracy and completeness.
@@ 5 Conclusion
[…]
@@
6 Evaluation on IMC dataset
Type Method IMC Dataset
AUC@3°$\degree$ AUC@5°$\degree$ AUC@10°$\degree$
Detector-Based COLMAP (SIFT+NN) 23.58 23.66 44.79
SIFT + NN + PixSfM 25.54 34.80 46.73
D2Net + NN + PixSfM 8.91 12.26 16.79
R2D2 + NN + PixSfM 31.41 41.80 54.65
SP + SG + PixSfM 45.19 57.22 70.47
Detector-Free LoFTR + PixSfM 44.06 56.16 69.61
LoFTR + DFSfM 46.55 58.74 72.19
AspanTrans. + DFSfM 46.79 59.01 72.50
MatchFormer + DFSfM 45.83 57.88 71.22
Dense Matching RoMa + DFSfM 47.43 59.84 73.19
Ours (DKM Quantized) 48.65 61.09 74.37
Ours (RoMa Quantized) 48.48 60.79 73.90
Deep VGGSfM 45.23 58.89 73.92
Table 4: Results of Multi-View Camera Pose Estimation on IMC Dataset. Our framework is compared with detector-based, detector-free and dense-matching baselines on the IMC Phototourism datasets by the AUC of pose error at different thresholds. Bold and underline indicate the best and second-best results.
Dataset.
For completeness, we report the performance of our method on the Image Matching Challenge (IMC) 2021 dataset [25]. The IMC Phototourism dataset comprises large-scale outdoor scenes with significant viewpoint, illumination variations, and temporal occlusions. To obtain dense matching from image pairs, we resize the longer edge to 1162 pixels, and the other edge to 768 pixels. Non-max suppression is applied with a radius of 2 pixels.
Evaluation.
Our evaluation includes all test scenes, with results reported at varying thresholds and averaged across all scenes. Following the evaluation protocol outlined in the main paper, we assess the accuracy of the estimated multi-view poses using the AUC of pose error. Due to the high occlusion and appearance changes in many test scenes, directly applying track extension via Gaussian Splatting is challenging. Instead, we adopt a quantization method for the matching results, focusing on comparing our multi-view kernelized matching module with DFSfM [24]’s refinement module. Specifically, we apply quantization to the matched keypoints with a radius of $4$, following the same protocol as DFSfM, and then use our refinement module to refine the quantized keypoints.
Results.
As shown in Tab. 4, our refinement module with dense matching outperforms existing baselines over all metrics. As exemplified in Fig. 7, our framework successfully reconstructs dense and accurate 3D points leading to improved pose estimation accuracy.
Figure 7: Qualitative Results on IMC dataset. Our method with dense matching RoMa is qualitatively compared with the baseline LoFTR + DFSfM. The red cameras are ground-truth poses while the blue cameras are recovered poses.
7 Ablative study
Refinement Camera Pose Estimation Accu.(%) (Pipes)
AUC@1°$\degree$ AUC@3°$\degree$ AUC@5°$\degree$ 1cm 2cm 5cm
No Refine 39.20 66.51 74.12 55.02 78.91 93.90
Iter 1 60.83 78.26 82.54 66.35 85.71 95.95
Track Extension via GS + Iter 2 60.92 78.41 82.63 74.75 92.06 97.98
Matching Design Scene Name Avg.
Average Track Length
Raw 2.19 2.06 2.12 2.41 2.09 2.08 2.07 2.03 2.04 2.08 2.05 2.12 2.08 2.11
As shown in Tab. 5, our pipeline improves the accuracy of both pose and point cloud through refinement. Note that the point cloud of No Refine is triangulated from two-view non-quantized matching results (resulting in most tracks having a length of 2), based on coarse poses obtained from quantized matching as DFSfM. Thanks to track extension via Gaussian Splatting, we can feed more views into multi-view kernelized refinement module, resulting in a slight improvement on camera pose accuracy, and a significant improvement on point cloud accuracy.
8 Details of Track extension via Gaussian Splatting
8.1 Training Gaussian
Initialization.
As described in Sec.3.2 in the main paper, 3D Gaussians are initialized based on the initial SfM point cloud. We set the initial Gaussians’ position to the 3D point coordinates, opacity to 1, and rotation to the identity matrix. The scale parameter $S$ is determined using the method in [26], where each Gaussian’s radius is set such that it is projected as an one-pixel radius circle on the 2D image. Specifically, we compute depth values for each 3D point based on the camera poses of its keypoints. Then we select the maximum value across the depth values and divide it by the focal length as follows:
$S=\frac{D_{max}}{f}$ (9)
Training details.
Unlike the standard training [27] that requires over 10,000 iterations for high-quality rendering, our process focuses only on verifying the visibility of the initial SfM’s 3D points. Thus, we do not need as many iterations as the general training for high quality image rendering [27, 75].
8.2 Track length analysis
9 Sparse feature matching with our refinement
@@
Our track refinement module (multi-view kernelized matching described in Sec.3.3) can also be applied to refine SfM models reconstructed using sparse detection and matching to further improve the robustness of SfM. We evaluate each SfM result on the ETH3D dataset, assessing accuracy and completeness following the ETH3D benchmark [53]. As shown in Tab. 7, our refinement module consistently outperforms PixSfM [30] and DFSfM in accuracy when accompanied by the same sparse detectors and matchers. Furthermore, our refinement module does not compromise the completeness much for higher accuracy.
10 Semi-dense matching with our pipeline
@@ 11 Applying Gaussian Splatting
To demonstrate the utility of our framework, we use it to initialize 3D Gaussians for novel view synthesis in a few-shot setting. Starting from SfM points, we apply Few-Shot Gaussian Splatting (FSGS)[75] on the LLFF dataset[35]. We compare the performance of 3D Gaussian models initialized using SIFT [31] with COLMAP, DetectorFree-SfM with LoFTR, and our framework with RoMa, using fixed camera poses and intrinsics provided by [75]. For Gaussian initialization and training, we use only three views and render the original image size provided in the dataset.
@@
12 Runtime comparison
We evaluate the runtime of our multi-view kernelized matching module against the multi-view matcher module in [24] on the Pipes scene, using the same SfM model for refinement. Our proposed refinement module costs a runtime increase about 19%, from 37.1s to 44.2s on a single RTX 4090 GPU, due to the added computational cost of the Gaussian Process and CNN Decoder.
@@ 13 Limitations and future works
A key limitation of our framework is that track extension via Gaussian Splatting struggles in scenes with high photometric variation or transient occlusions. As future work, our pipeline can be extended with more advanced Gaussian Splatting models, such as [28, 72, 6, 69], to better handle occlusions and appearance changes, allowing our framework to perform robustly in a wider range of scenarios.

### v2

@@ 4.1 Datasets
We use the ETH3D [54], Texture-Poor SfM datasets [25], and Image Matching Challenge (IMC) 2021 dataset [26], each presenting unique challenges across different types of scenes. The ETH3D dataset includes indoor and outdoor scenes with sparsely captured, high-resolution images, and ground truth poses calibrated using Lidar. The Texture-Poor SfM dataset focuses on object-centric scenes with low texture, comprising subsampled image sets. We use all test scenes from the IMC Phototourism dataset, evaluating them with subsampled image bags. The main challenge of this dataset lies in its sparse views, accompanied by significant variations in viewpoint and illumination.
@@ 4.2 Implementation Details
For efficiency, we apply non-max suppression [9] with a radius of 4 pixels for ETH3D and 2 pixels for Texture-Poor SfM and IMC dataset. For mutual verification, we filter out matches with a distance exceeding $\epsilon_{\textit{p}}=3$px in Eq. 2.
To determine the visibility of Gaussians (SfM points) from a camera pose, we set the visibility threshold $\epsilon_{\textit{v}}$ to 0.5 in Eq. 3. Details of training and implementation of Gaussian Splatting are provided in the supplementary material.
We use S2Dnet [19] as a CNN backbone for our refinement architecture, and adopt the multi-view transformer architecture from DFSfM [25]. To ensure a fair comparison, we apply the same refinement parameters as in DFSfM, including a local patch size ($p=15$) from the feature maps ($\mathbf{F}_{R}$ and $\mathbf{F}_{Q}$) obtained from the CNN backbone, and a search boundary for reference view’s refined keypoints set to $w=7$. The number of distribution anchors $C^{2}$ is set to $7^{2}$. For the CNN decoder in Fig. 4, we use a series of residual blocks [23] and channel attention blocks [73]. During bundle adjustment, matches with reprojection errors exceeding $\epsilon_{\textit{f}}=3$px are discarded. The total SfM refinement process is iterated twice. More implementation details are reported in the supplementary material.
4.3 3D Triangulation
Evaluation.
We use the ETH3D training set, comprising 13 indoor and outdoor scenes with millimeter-accurate dense point clouds as ground truth. Following the protocol in PixSfM [32] and DFSfM [25], we triangulate 3D points using fixed camera poses and intrinsics, then evaluate the resulting point clouds for accuracy and completeness using the ETH3D benchmark [54]. The results are reported at varying distance thresholds (1cm, 2cm, 5cm), averaged across all scenes. Consistent with the benchmark setup, accuracy and completeness are measured by the proportion of triangulated points that fall within the specified distances from the ground truth points.
@@ Results.
The results are shown in Tab. 1. While there is a trade-off between accuracy and completeness, our Dense-SfM framework achieves superior performances on both metrics. Compared to the state-of-the-art baseline LoFTR + DFSfM, our pipeline with RoMa achieves higher accuracy and completeness, thanks to employing dense matching and our proposed pipeline. Also, compared to the RoMa + DFSfM, where we apply quantization on dense matching and refine the tracks as in DFSfM, our method achieves higher performance in terms of both accuracy and completeness, highlighting the effectiveness of our SfM framework. More results with detector-based and semi-dense matching are reported in the supplementary material.
Type Method ETH3D Dataset Texture-Poor SfM Dataset IMC Dataset
AUC@1°$\degree$ AUC@3°$\degree$ AUC@5°$\degree$ AUC@3°$\degree$ AUC@5°$\degree$ AUC@10°$\degree$ AUC@3°$\degree$ AUC@5°$\degree$ AUC@10°$\degree$
Detector-Based COLMAP (SIFT+NN) [52] 26.71 38.86 42.14 2.87 3.85 4.95 23.58 23.66 44.79
SIFT [33] + NN + PixSfM [32] 26.94 39.01 42.19 3.13 4.08 5.09 25.54 34.80 46.73
D2Net [12] + NN + PixSfM 34.50 49.77 53.58 1.54 2.63 4.54 8.91 12.26 16.79
R2D2 [45] + NN + PixSfM 43.58 62.09 66.89 3.79 5.51 7.84 31.41 41.80 54.65
SP [9] + SG [50] + PixSfM 50.82 68.52 72.86 14.00 19.23 24.55 45.19 57.22 70.47
Semi-Dense Matching LoFTR [59] + PixSfM 54.35 73.97 78.86 20.66 30.49 42.01 44.06 56.16 69.61
LoFTR + DFSfM [25] 59.12 75.59 79.53 26.07 35.77 45.43 46.55 58.74 72.19
AspanTrans. [6] + DFSfM 57.23 73.71 77.70 25.78 35.69 45.64 46.79 59.01 72.50
MatchFormer [69] + DFSfM 56.70 73.00 76.84 26.90 37.57 48.55 45.83 57.88 71.22
Dense Matching RoMa + DFSfM 57.97 77.02 81.59 34.28 50.01 69.01 47.43 59.84 73.19
DKM + Ours 59.04 77.73 82.20 41.19 57.26 72.95 48.65 61.09 74.37
RoMa + Ours 60.92 78.41 82.63 49.94 66.23 81.41 48.48 60.79 73.90
Deep VGGSfM [67] 28.09 44.41 49.91 1.86 6.26 21.10 45.23 58.89 73.92
Point-Based Mast3r-SfM [11] 35.85 58.46 65.03 43.46 61.51 78.97 42.26 54.53 67.97
Mast3r-SfM + Our Refinement 37.50 59.18 65.48 48.03 64.48 80.31 44.98 57.09 70.09
@@ 4.4 Multi-View Camera Pose Estimation
For the evaluation, we use a subset of ETH3D training and test dataset, which includes 22 indoor and outdoor scenes, the Texture-Poor SfM dataset, and the test scenes of IMC 2021 dataset. We evaluate the accuracy of the estimated multi-view poses using the AUC of pose error at various thresholds. For the evaluation of IMC 2021 dataset, due to the high occlusion and appearance changes in many test scenes, directly applying track extension via GS is challenging. Instead, we adopt a quantization method [25] on the matching results, focusing on comparing our multi-view kernelized matching module with DFSfM’s refinement module. Specifically, we apply quantization to the matched keypoints with a radius of $4$, and then use our refinement module to refine the quantized keypoints.
Figure 6: Qualitative Results on Texture-Poor SfM and IMC dataset. Our method with the dense matcher RoMa is qualitatively compared with the baseline LoFTR + DFSfM. The red cameras are ground truth poses while the blue cameras are recovered poses.
@@ 4.5 Ablation Studies
We conduct several experiments to evaluate the key components of our pipeline, including the matching design used as input for SfM and the architectural choices of the refinement module, on the ETH3D dataset using triangulation metrics.
@@
For a comparison with DFSfM [25], which applies quantization to matched keypoints to address the fragmented track problem, we replaced our GS track extension with their quantization approach, using a rounding ratio of $r=4$. Results indicate that our GS track extension not only maintains completeness by avoiding the limitations of quantization but also achieves superior accuracy.
@@
We first compare our refinement module to the multi-view refinement module in DFSfM [25], by substituting our refinement module with it. The result shows that our refinement module consistently outperforms the existing approach, achieving higher accuracy and completeness.
@@ 5 Conclusion
6 Acknowledgement
This research was supported in part by Samsung Advanced Institute of Technology (SAIT), the NAVER-Intel Co-Lab, Inter-university Semiconductor Research Center (ISRC), and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)].
@@ 6 Acknowledgement
7 Experiment
7.1 Sparse feature matching with our refinement
@@
Our track refinement module (multi-view kernelized matching described in Sec.3.3) can also be applied to refine SfM models reconstructed using sparse detection and matching to further improve the robustness of SfM. We evaluate each SfM result on the ETH3D dataset, assessing accuracy and completeness following the ETH3D benchmark [54].
As shown in Tab. 4, our refinement module consistently outperforms PixSfM and DFSfM in accuracy when accompanied by the same sparse detectors and matchers. Furthermore, our refinement module does not compromise the completeness much for higher accuracy.
7.2 Semi-dense matching with our pipeline
@@
Matching Design Scene Name Avg.
Average Track Length
Raw 2.19 2.06 2.12 2.41 2.09 2.08 2.07 2.03 2.04 2.08 2.05 2.12 2.08 2.11
@@
7.3 Implementation Details
As in PixSfM and DFSfM, we obtain match results from exhausitive pairs within a set of images. To obtain dense matching from image pairs, we resize images to $1162\times 768$ for the ETH3D dataset and IMC 2021 dataset, while maintaining the original size of $840\times 840$ for the Texture-Poor SfM dataset.
Since dense matching methods (e.g. DKM and RoMa) are computationally heavier than detector-based or semi-dense approaches due to their pixel-wise matching process, we improve efficiency by obtaining coarse matches at a lower resolution ($560\times 560$ for the ETH3D and IMC 2021, $420\times 420$ for the Texture-Poor SfM dataset) in the global matching stage of RoMa, which dominates the runtime. We then refine the matches to the target resolution. Under this configuration, on the ETH3D dataset, RoMa achieves a runtime of 0.08 seconds per image pair on a single A6000 GPU. In comparison, DFSfM with LoFTR operates at a higher resolution ($1600\times 1200$) and requires 0.2 seconds per pair. Our matching pipeline, which includes bidirectional matching with verification, runs in 0.17 seconds per pair, outperforming the LoFTR-based pipeline while maintaining a comparable matching speed.
7.4 More Ablation Study
Refinement Camera Pose Estimation Accu.(%) (Pipes)
AUC@1°$\degree$ AUC@3°$\degree$ AUC@5°$\degree$ 1cm 2cm 5cm
No Refine 39.20 66.51 74.12 55.02 78.91 93.90
Iter 1 60.83 78.26 82.54 66.35 85.71 95.95
Track Extension via GS + Iter 2 60.92 78.41 82.63 74.75 92.06 97.98
As shown in Tab. 7, our pipeline improves the accuracy of both pose and point cloud through refinement. Note that the point cloud of No Refine is triangulated from two-view non-quantized matching results (resulting in most tracks having a length of 2), based on coarse poses obtained from quantized matching as DFSfM. Thanks to track extension via Gaussian Splatting, we can feed more views into our multi-view kernelized refinement module, resulting in a slight improvement on camera pose accuracy, and a significant improvement on point cloud accuracy.
7.5 Runtime Comparison
We evaluate the runtime of our multi-view kernelized matching module against the multi-view matcher module in DFSfM[25] on the Pipes scene, using the same SfM model for refinement. Our proposed refinement module costs a runtime increase about 22%, from 35.3s to 43.2s on a single A6000 GPU, due to the added computational cost of the Gaussian Process and CNN Decoder.
8 Details of Track extension via Gaussian Splatting
8.1 Training Gaussians
Initialization.
As described in Sec.3.2 in the main paper, 3D Gaussians are initialized based on the initial SfM point cloud. We set the initial Gaussians’ position to the 3D point coordinates, opacity to 1, and rotation to the identity matrix. The scale parameter $S$ is determined using the method in Splatem [27], where each Gaussian’s radius is set such that it is projected as an one-pixel radius circle on the 2D image. Specifically, we compute depth values for each 3D point based on the camera poses of its keypoints. Then we select the maximum value across the depth values and divide it by the focal length as follows:
$S=\frac{D_{max}}{f}$ (9)
Training details.
Unlike the standard training of GS [28] that requires over 10,000 iterations for high-quality rendering, our process focuses only on verifying the visibility of the initial SfM’s 3D points. Thus, we do not need as many iterations as the general training for high quality image rendering [28, 77].
8.2 Track length analysis
@@ 9 Applying Gaussian Splatting
To demonstrate the utility of our framework, we use it to initialize 3D Gaussians for novel view synthesis in a few-shot setting. Starting from SfM points, we apply Few-Shot Gaussian Splatting (FSGS) [77] on the LLFF dataset[36]. We compare the performance of 3D Gaussian models initialized using SIFT [33] with COLMAP, DetectorFree-SfM (DFSfM) with LoFTR, and our framework with RoMa, using fixed camera poses and intrinsics provided by FSGS [77] pipeline. For Gaussian initialization and training, we use only three views and render the original image size provided in the dataset.
@@ 10 Limitations and Future works
A key limitation of our framework is that track extension via Gaussian Splatting struggles in scenes with high photometric variation or transient occlusions. As future work, our pipeline can be extended with more advanced Gaussian Splatting models [29, 74, 7, 71], to better handle occlusions and appearance changes, allowing our framework to perform robustly in a wider range of scenarios.